# Asynchronous-Many-Task Systems: Challenges and Opportunities - Scaling an AMR Astrophysics Code on Exascale machines using Kokkos and HPX

### v1

@@ 1 Introduction
Adaptive mesh refinement (AMR) is a crucial component for many high-resolution, multi-physics, and multi-model simulations. In this work, we are focusing on one such simulation code: the astrophysics application Octo-Tiger. Octo-Tiger simulates binary star systems for studying their interactions and the various phenomena caused by them. In these simulations, large computational domains are needed, for example, to track the mass leaving the stars; see the streamlines in Figure 1. To resolve the quantity of interest, such as the concentration of $^{16}\textrm{O}$, some areas of the large computational domain, like the stars (donor and accretor) and accretion column, require a high resolution, which is why AMR is such a crucial component. To provide an example of such an adaptively refined mesh with Octo-Tiger, Figure 2 shows the mesh close to the merger on the equatorial plane. The color shows the portion of the grid assigned to various computational nodes. The workload of the nodes is unbalanced concerning the domain, resulting in differing sizes of the allocated domains to the nodes to achieve load balancing across the mesh. In addition, some nodes have more cells due to the refined mesh; where the smallest cells have a size of $1.4\text{\times}{10}^{7}$ cm and the largest cells have a dimension of $2.24\text{\times}{10}^{8}$ cm.
@@
Given how widespread these concerns are for developing scalable AMR applications there are existing solutions available to address them. Specifically, for our work, we turn to HPX and Kokkos.
@@
HPX, a distributed asynchronous task-based runtime system, can help us to tackle scalability. Expressing the parallelism within Octo-Tiger’s tree-structure with the dynamic HPX task-graph and utilizing HPX’s distributed features helps with many of the aforementioned challenges, for instance easing the development effort for the distributed memory management, transparent overlapping of computation and communictation, and exploiting the parallelism during Octo-Tiger’s solver iterations. Kokkos, a framework for developing portable compute kernels, in turn helps us with portability in terms of code and performance.
Thus, in theory, combining HPX and Kokkos presents an opportunity to develop distributed and portable HPC applications more easily, helping to achieve scalability on a wide range of CPU and GPU supercomputers.
However, in practice, we have found that this approach required some additional glue in order for the frameworks to work together seamlessly when developing an AMR application, such as Octo-Tiger. Hence, we needed to develop some integrations and add missing pieces and tools where necessary.
@@
In some scenarios, the frameworks themselves did not work together well (such as Kokkos’ internal fences (barriers) needlessly blocking HPX worker threads instead of suspending the HPX task). Also we encountered problems regarding missing tools to adapt our use-case to the given hardware. For example, Octo-Tiger was originally developed only for CPUs. When porting Octo-Tiger to Kokkos, the workload per tree-node proved too small to provide sufficient work for an efficient compute kernel (especially on GPUs), causing poor performance due to GPU device starvation. This is a problem shared by many tree-based codes. Moreover, the dynamic nature of the work (with tree-nodes being added, removed, and migrated over the course of the simulation) greatly complicates any static approach for aggregating the workload of different tree-nodes to address this.
@@
Over the recent years, we faced and addressed many of these issues and missing pieces in our previous work. Notably, we aimed to do so in a way that makes our solutions usable universally in codes other than Octo-Tiger: We improved the interoperability of HPX and Kokkos considerably, allowing us to directly and asynchronously integrate Kokkos operations into the HPX task-graph, eliminating the need to block CPU threads with blocking fence operations entirely [1]. We addressed the device starvation issue by implementing a set of allocators and executors that facilitate dynamic work aggregation, combining individual kernel launches of compatible kernels on-the-fly into a single large compute kernel guided by the current GPU work load [2]. Furthermore, we made use of SIMD types within our Kokkos kernels (allowing for both the Kokkos SIMD types and the std::simd types), to improve the efficiency on CPU platforms whilst seamlessly retaining GPU support. Here, we contributed types for SVE [3] and RVV [4] to improve platform support, especially for Fugaku [5]. We also added an improved networking backend to HPX [6]. Each of these previous works considerably improved the usability of HPX and Kokkos for developing scalable and portable AMR applications, thus helping us to realize the opportunities that the combined usage of HPX and Kokkos offer. Over the course of this previous work, we also ported Octo-Tiger’s most important compute kernels to Kokkos, turning it into an GPU-accelerated application.
@@
In this work, we build on these achievements, combine all methodologies, enable the seamless exchange of different backends on cutting-edge systems, and show novel scalability results up to full system runs for three of the world’s fastest supercomputers with entirely different hardware. The presented work thus builds directly on our previous work, combining those results in Octo-Tiger to enable running it on multiple major supercomputers, examining the scalability and performance of production scenarios at scale on these systems.
@@
Additionally, in this work, we provide a better overview of all of our previous results to make HPX and Kokkos more suitable for applications such as Octo-Tiger. This is to provide the reader with a better insight into what features and framework integrations we had to implement to scale our AMR application to GPU supercomputers in the first place. Crucially, these features and integrations all work independent of Octo-Tiger and are available for other HPC application developers in case they consider using HPX and Kokkos. Additionally, this overview may prove to be useful for developers of frameworks similar to HPX: it outlines what challenges need to be overcome for efficient use of a distributed task-based framework with Kokkos, especially when dealing with GPUs and tree-based user codes.
The remainder of this work is structured as follows: First, we will mention related work in the next section. Then, as Octo-Tiger is a major part of this work, serving both as motivation and as a benchmark for many of our developments, we will cover its astrophysical use-case and the computational challenges in greater detail afterwards. Next, we will introduce HPX, Kokkos, and Octo-Tiger themselves. Then, we will provide the overview of our previous work that addressed the challenges we found when using HPX and Kokkos in a distributed AMR application. After that, we move to describing the distributed runs we performed on various machines with Octo-Tiger, realizing the opportunities presented by cobining HPX and Kokkos, namely scalability and portability. Lastly, given the runtime data collected in the previous section, we discuss our next steps to improve Octo-Tiger and conclude this work.
@@ 2 Related Work
Concerning the astrophysics application, the CASTRO [7] application supports adaptive mesh refinement with similar astrophysics features. Castro uses the MPI+X approach where X is OpenMP for CPUs and X is either CUDA for NVIDIA GPUs or RocM for AMD GPUs. Castro uses AMReX [8] for its adaptive mesh refinement. From the code portability perspective, the following alternatives to Kokkos are available: OpenACC, OpenCL, Raja [9], and SYCL. We decided to use Kokkos for the two following reasons: 1) we wanted to use a C++ library since HPX leverages the C++ standard and that would simplify the integration and 2) AMD, NVIDIA, and Intel GPU support was a strong requirement. For the GPUs, we use Kokkos GPU execution spaces. For the kernels on the CPU, we use the Kokkos execution space Kokkos::HPX [1] to execute the compute kernel on HPX’s light-weight threads. In addition, we use HPX-Kokkos to launch Kokkos’ API functions asynchronously and integrate those within HPX’s task graph. HPX is not the only asynchronous many-task system (AMT) with GPU capabilities in existance. Since this work focuses on distributed runs, we will mention only distributed AMTs. Table 1 shows the GPU support of Chapel [10], Charm++ [11], Legion [12], Uintah [13], and PaRSEC [14]. For a detailed survey, we refer to [15]. Only Legion and Uintah have Kokkos support. Uintah adopts Kokkos through a Uintah-specific intermediate portability layer used to preserve legacy code and centralize Kokkos calls [16]. Legion allows for Kokkos API calls. To conclude, the novelty of our work is that Legion and Uintah solely use Kokko’s GPU execution space while we additionally use Kokkos’ HPX backend to launch compute kernels on the CPU.
@@ 3.1 The Astrophysical Problem
The majority of stars in the Universe are not isolated stars but, rather, members of gravitationally bound systems of stars with two or more components. Binary stars are star systems with two components that form from the same cloud of gas. The components of binary systems eventually may evolve to the “white dwarf” stage, where nuclear burning has ceased, much of the stellar material has been ejected, and remaining remnants emit only black-body radiation. Known as a double white dwarf (DWD), the components may be driven closer together through emission of gravitational radiation. When the components of a DWD are close enough that the gravity of the “accretor” pulls mass from the “donor,” this is said to be an “interacting” DWD (see Figure 1). Interacting binary systems are commonplace in the Universe. A large fraction of interacting binaries result in the disruption of the donor and subsequent merger of much of its material with the accretor, which makes understanding them crucial to our understanding of a wide variety of potential outcomes for such mergers (e.g., type 1A super-novae, R Coronae Borealis stars).
The R Coronae Borealis stars have strange elemental abundances, being low in hydrogen but high in carbon content. They are variable stars, fading from and then returning to maximum brightness at intervals of several years. This is thought to be caused by clouds of carbon dust surrounding the star. They are thought to form from the merger of double white dwarfs. Octo-Tiger has been used extensively in the past to study this question. e.g.  [17, 18, 19, 20]. Octo-Tiger has also been used to investigate whether the star Belelgeuse may be the outcome of a past merger  [21].
The numerical modelling of the short-lived phase of unstable mass transfer in a DWD merger requires a fully three-dimensional, self-consistent treatment of the effects of hydrodynamics and gravity in a rotating frame of reference. There are two basic approaches to the hydrodynamics: Smoothed Particle Hydrodynamics (SPH) and the grid-based finite volume method (FVM). One of the processes crucial to understanding the formation of R Coronae Borealis stars is the creation of an unusually high concentration of $^{16}\textrm{O}$ in the accretor during merger. SPH codes tend to greatly underestimate this effect, making it difficult to resolve the low mass transfer rates typical of most of the merger processes. FVM codes tend to overestimate it. With FVM, however, it is possible to conduct convergence studies by successively doubling grid resolution to better understand the magnitude of the overestimation.
@@ 3.2 The computational challenges
The need for application codes to be portable—both in terms of code and performance—is one of the central problems of modern HPC computing. The quickly growing complexity of hardware architectures and software stacks require rapid advancements in development productivity, performance and software portability. This work relies on standard C++ and widely used portable C++ libraries, such as Kokkos (a C++ Performance Portability Programming Ecosystem [22, 23]), HPX (a fully C++ standards conforming asynchronous many-task runtime system [24, 25, 26]), and C++ standard SIMD features that have been added to C++26 [27] to help ensure full code portability of our application. The use of Kokkos, HPX, and std::simd guarantees its seamless portability and consistent scaling characteristics across a wide variety of heterogeneous (GPU and non-GPU based) hardware platforms, in particular on Intel, AMD, and NVIDIA products, but also on upcoming ARM64 and RISC-V HPC architectures.
In the problem domain investigated, the amount of computational work per compute kernel is low, requiring adaptive work aggregation on GPU devices to reduce scheduling overheads and starvation effects for the related asynchronous tasks [2]. The nature of highly complex 3D AMR applications like Octo-Tiger (see Figure 2) require Exa-Scale computing systems to solve underlying physics problems that require very large amounts of data and pose computational challenges. To achieve high parallel efficiency, these challenges include providing intra- and inter-node load-balancing, pervasive overlapping of computation and communication, tight integration and high utilization of accelerator devices, and runtime-adaptive task placement and scheduling across nodes and diverse execution environments.
@@ 4.1 HPX
HPX is an asynchronous many-task runtime system (AMT). It allows users to express the parallelism within their application by dynamically building a task-graph on-the-fly using HPX futures and continuations. For example, a function can be launched asynchronously using hpx::async which returns an HPX future. This future can be used to suspend the current HPX task (get), or define another task that will be automatically triggered once the original function is done (then). Each task can subsequently spawn an arbitrary number of other tasks asynchronously and with little overhead, making parallel work available quickly, which is especially useful for parallelized tree traversals. Of course, there is a lot more functionality available. Particularly, HPX implements all C++23 APIs regarding parallelism and concurrency. For a full list of the functionality available, we refer to HPX’s online documentation [^1] https://hpx-docs.stellar-group.org/latest/html/index.html. The tasks within HPX’s task graph are then processed by a set of HPX worker threads. Thus, HPX can easily handle billions of tasks, which are processed by the worker threads as they become available within the graph.
@@
HPX also contains various key features to enable the development of distributed codes [28]. For example, functions can also be asynchronously launched on other compute-nodes, still getting an HPX future back. There are also communication channels and other distributed features available that tie into the HPX task graph similarly. Essentially, this allows users to build a distributed task graph. This is enabled by HPX’s Active Global Address Space (AGAS) and various networking backends (parcelports).
@@
Overall, HPX is well suited to address the computational challenges of developing scalable, distributed AMR applications (directly addressing some of the challenges mentioned in Section The computational challenges). With it, users can exploit the parallelism within tree-structures more easily by expressing it in a similar graph structure, the HPX task-graph. Furthermore, through this task graph (which may contain both computation and communication tasks), users can automatically achieve an overlapping of computation and communication, helping them to scale their codes to numerous compute nodes with a high parallel efficiency. Furthermore, thanks to some of our integrations, we can also incorporate accelerator devices into this task-graph, helping to keep a high utilization of the accelerator devices by better overlapping their computations with the CPU tasks and communication.
@@
Additionally, using HPX can also help with portability. HPX’s portability is achieved through an interface (API) fully aligned with the latest C++ standard specification. The API decouples the application from the upper layers of the runtime system and hides the low-level details of the underlying architecture. HPX provides general-purpose building blocks (such as C++ standards conforming parallel algorithms), as well as high-level utilities coordinating asynchronous execution, and offers a programming abstraction that facilitates the programming of both shared-memory and distributed-memory systems through a uniform programming interface, scaling up from hand-held devices to HPC clusters.
@@
In terms of networking portability, HPX’ modular architecture abstracts the concrete networking hardware allowing for Octo-Tiger to be completely independent of it. HPX supports TCP/IP, MPI, LCI [6, 29], OpenSHMEM [30], and GASNet [31] as possible networking conduits. In this work, we focus on two backends (parcelports): MPI and LCI.

### v2

@@ 1 Introduction
Adaptive mesh refinement (AMR) is a crucial component for many high-resolution, multi-physics, and multi-model simulations. In this work, we are focusing on one such simulation code: the astrophysics application Octo-Tiger. Octo-Tiger simulates binary star systems, allowing one to study their interactions and the various phenomena caused by them. In these simulations, large computational domains are needed, for example, to track the mass leaving the stars; see the streamlines in Figure 1. To resolve the quantity of interest, such as the concentration of $^{16}\textrm{O}$, some areas of the large computational domain, like the stars (donor and accretor) and accretion column, require a high resolution, which is why AMR is such a crucial component. To provide an example of such an adaptively refined mesh with Octo-Tiger, Figure 2 shows the mesh close to the merger on the equatorial plane. The color shows the portion of the grid assigned to various computational nodes. The workload of the nodes is unbalanced with respect to the domain. This results in domains of differing sizes being allocated across nodes to achieve load balancing across the mesh. In addition, some nodes have more cells due to the refined mesh; where the smallest cells have a size of $1.4\text{\times}{10}^{7}$ cm and the largest cells have a dimension of $2.24\text{\times}{10}^{8}$ cm.
@@
Given how widespread these concerns are for developing scalable AMR applications, there are existing solutions available to address them. Specifically, for our work, we turn to HPX and Kokkos. HPX, a distributed asynchronous task-based runtime system, can help us to tackle scalability. Expressing the parallelism within Octo-Tiger’s tree-structure with the dynamic HPX task-graph and utilizing HPX’s distributed features helps with many of the aforementioned challenges, for instance easing the development effort for the distributed memory management, transparent overlapping of computation and communication, and exploiting the parallelism during Octo-Tiger’s solver iterations. Kokkos, a framework for developing portable compute kernels, in turn helps us with portability in terms of code and performance. Thus, in theory, combining HPX and Kokkos presents an opportunity to develop distributed and portable HPC applications more easily, helping to achieve scalability on a wide range of CPU and GPU supercomputers.
@@
However, in practice, we have found that this approach required some additional glue in order for the frameworks to work together seamlessly when developing an AMR application, such as Octo-Tiger. Hence, we needed to develop some integrations and add missing pieces and tools where necessary. In some scenarios, the frameworks themselves did not work well together (such as Kokkos’ internal fences (barriers) needlessly blocking HPX worker threads instead of suspending the HPX task). We also encountered problems regarding missing tools to adapt our use-case to the given hardware. For example, Octo-Tiger was originally developed only for CPUs. When porting Octo-Tiger to Kokkos, the workload per tree-node proved too small to provide sufficient work for an efficient compute kernel (especially on GPUs), causing poor performance due to GPU device starvation. This is a problem shared by many tree-based codes. Moreover, the dynamic nature of the work (i.e., tree-nodes being added, removed, and migrated over the course of the simulation) greatly complicates any static approach for aggregating the workload of different tree-nodes to address this.
@@
Over the recent years, we faced and addressed many of these issues and missing pieces in our previous work. Notably, we aimed to do so in a way that makes our solutions usable universally in codes other than Octo-Tiger. As a part of past work, we improved the interoperability of HPX and Kokkos considerably, allowing us to directly and asynchronously integrate Kokkos operations into the HPX task-graph, eliminating the need to block CPU threads with blocking fence operations entirely [1]. We addressed the device starvation issue by implementing a set of allocators and executors that facilitate dynamic work aggregation, combining individual kernel launches of compatible kernels on-the-fly into a single large compute kernel guided by the current GPU work load [2]. Furthermore, we made use of SIMD types within our Kokkos kernels (allowing for both the Kokkos SIMD types and the std::simd types), to improve the efficiency on CPU platforms whilst seamlessly retaining GPU support. Here, we contributed std::simd-compatible SIMD types for Scalable Vector Extension (SVE) [3] and RISC-V Vector Extension (RVV) [4] to improve platform support, especially for Fugaku [5]. We also added an improved networking backend to HPX [6]. Each of these previous works considerably improved the usability of HPX and Kokkos for developing scalable and portable AMR applications, thus helping us to realize the opportunities that the combined usage of HPX and Kokkos offers. Over the course of this previous work, we also ported Octo-Tiger’s most important compute kernels to Kokkos, turning it into a GPU-accelerated application.
@@
In this work, we build on these achievements, combine all methodologies, enable the seamless exchange of different backends on cutting-edge systems, and show novel scalability results up to full system runs for three of the world’s fastest supercomputers with entirely different hardware. Thus, the work presented here builds directly on our previous work, combining those results in Octo-Tiger to allow one to run it on multiple major supercomputers and examine the scalability and performance of production scenarios at scale on these systems.
@@
Additionally, in this work, we provide a better overview of all of our previous results to make HPX and Kokkos more suitable for applications such as Octo-Tiger. This is to provide the reader with a better insight into what features and framework integrations we had to implement to scale our AMR application to GPU supercomputers in the first place. Crucially, these features and integrations all work independent of Octo-Tiger and are available for other HPC application developers in case they consider using HPX and Kokkos. As a result, this overview may also prove to be useful for developers of frameworks similar to HPX: it outlines what challenges need to be overcome for efficient use of a distributed task-based framework with Kokkos, especially when dealing with GPUs and tree-based user codes.
The remainder of this work is structured as follows: First, we will mention related work in the next section. Then, as Octo-Tiger is a major part of this work, serving both as motivation and as a benchmark for many of our developments, we will cover its astrophysical use-case and the computational challenges in greater detail afterwards. Next, we will introduce HPX, Kokkos, and Octo-Tiger themselves. Then, we will provide the overview of our previous work that addressed the challenges we found when using HPX and Kokkos in a distributed AMR application. After that, we move to describing the distributed runs we performed on various machines with Octo-Tiger, realizing the opportunities presented by combining HPX and Kokkos, namely scalability and portability. Lastly, given the runtime data collected in the previous section, we discuss our next steps to improve Octo-Tiger and conclude this work.
@@ 2 Related Work
Concerning the astrophysics application, the CASTRO [7] application supports adaptive mesh refinement using AMReX [8] with similar astrophysics features. Castro uses the MPI+X approach where X is OpenMP for CPUs and X is either CUDA for NVIDIA GPUs or HIP for AMD GPUs. Octo-Tiger and CASTRO are both astrophysical simulation codes, but they differ in numerical approach and intended use cases. CASTRO is a general-purpose, block-structured AMR code designed for a broad range of compressible flow problems, employing finite volume hydrodynamics and multigrid gravity solvers. Octo-Tiger, in contrast, is specialized for modeling rotating binary star systems. It uses an octree-based AMR mesh with finite volume hydrodynamics that uses the piecewise parabolic method (PPM) to compute the values at all 26 spatial directions on a grid cell. It includes a Self-Consistent Field (SCF) solver capable of generating synchronously rotating binary systems in equilibrium. This is a class of initial conditions that CASTRO cannot natively produce. Octo-Tiger also employs a Fast Multipole Method (FMM) for gravity, including the time derivative of the potential to ensure machine-precision energy conservation in the rotating frame. These differences in both software and scientific focus mean that direct performance comparisons between the two codes are not meaningful, as they are solving fundamentally different problems. Table 1 showcases the supported physics modules. Castro also provides nuclear reaction networks and thermal diffusion, which are not yet available in Octo-Tiger. Other MPI-based solutions are Paramesh [9], SAMRAI [10], FLASH [11], p4est [12], Enzo [13], Chombo [14], and deal.II [15]. Focusing on oct-tree based adaptive mesh refinement, tree traversals are needed and depending on the level of refinement $d$ this takes $\mathcal{O}(d)$ global reductions. For large scale simulations and high level of refinements, synchronization and communication overheads can affect the efficiency [16]. Potential solutions are 1) to use non-blocking MPI calls, 2) overlap communication and communication, and 3) load balancing. The asynchronous launches of HPX tasks easily facilitate non-blocking calls, describing tasks dependencies using HPX’s task graph allows for a natural overlap of communication and computation, and HPX’s support for work stealing of tasks between threads within a node helps with load balancing. A similar approach has been done using Charm++ [16].
Method Castro Octo-Tiger
Radiation
Gravity
Hydro
Nuclear reaction networks
Thermal diffusion
Table 1: Comparison of the physics modules available in Octo-Tiger and Castro.
From the code portability perspective, the following alternatives to Kokkos are available: OCCA, OpenACC, OpenCL, RAJA [17], and SYCL. We decided to use Kokkos for the two following reasons: 1) we wanted to use a C++ library since HPX leverages the C++ standard and that would simplify the development of our integrations and 2) AMD, NVIDIA, and Intel GPU support were strong requirements. For the GPUs, we use the respective Kokkos GPU execution spaces. HPX is not the only asynchronous many-task system (AMT) with GPU capabilities in existence. Since this work focuses on distributed runs, we will mention only distributed AMTs. Table 2 shows the GPU support of Chapel [18], Charm++ [19], Legion [20], Uintah [21], and PaRSEC [22]. For a detailed survey, we refer to [23]. Only Legion and Uintah have Kokkos support. Uintah adopts Kokkos through a Uintah-specific intermediate portability layer used to preserve legacy code and centralize Kokkos calls [24]. Legion allows for Kokkos API calls. To conclude, the novelty of our work is that Legion and Uintah solely use Kokkos’ GPU execution space while we additionally use Kokkos’ HPX backend to launch compute kernels on the CPU.
@@ 3.1 The Astrophysical Problem
The majority of stars in the Universe are not isolated stars but, rather, members of gravitationally bound systems of stars with two or more components. Binary stars are star systems with two components that form from the same cloud of gas. The components of binary systems may eventually evolve to the “white dwarf” stage, where nuclear burning has ceased, much of the stellar material has been ejected, and remaining remnants emit only black-body radiation. Known as a double white dwarf (DWD), the components may be driven closer together through emission of gravitational radiation. When the components of a DWD are close enough that the gravity of the “accretor” pulls mass from the “donor,” this is said to be an “interacting” DWD (see Figure 1). Interacting binary systems are commonplace in the Universe. A large fraction of interacting binaries result in the disruption of the donor and subsequent merger of much of its material with the accretor, which makes understanding them crucial to our understanding of a wide variety of potential outcomes for such mergers (e.g., type 1A super-novae, R Coronae Borealis stars).
The R Coronae Borealis stars have strange elemental abundances as they are low in hydrogen but high in carbon content. They are variable stars, fading from and then returning to maximum brightness at intervals of several years. This is thought to be caused by clouds of carbon dust surrounding the star. They are thought to form from the merger of double white dwarfs. Octo-Tiger has been used extensively in the past to study this question. e.g.  [25, 26, 27, 28]. Octo-Tiger has also been used to investigate whether the star Betelgeuse may be the outcome of a past merger  [29].
The numerical modeling of the short-lived phase of unstable mass transfer in a DWD merger requires a fully three-dimensional, self-consistent treatment of the effects of hydrodynamics and gravity in a rotating frame of reference. There are two basic approaches to the hydrodynamics: Smoothed Particle Hydrodynamics (SPH) and the grid-based finite volume method (FVM). One of the processes crucial to understanding the formation of R Coronae Borealis stars is the creation of an unusually high concentration of $^{16}\textrm{O}$ in the accretor during merger. SPH codes tend to greatly underestimate this effect, making it difficult to resolve the low mass transfer rates typical of most of the merger processes. FVM codes tend to overestimate it. With FVM, however, it is possible to conduct convergence studies by successively doubling grid resolution to better understand the magnitude of the overestimation.
@@ 3.2 The computational challenges
The need for application codes to be portable—both in terms of code and performance—is one of the central problems of modern HPC computing. The quickly growing complexity of hardware architectures and software stacks require rapid advancements in development productivity, performance, and software portability. This work relies on standard C++ and widely used portable C++ libraries, such as Kokkos (a C++ Performance Portability Programming Ecosystem [30, 31]), HPX (a fully C++ standards conforming asynchronous many-task runtime system [32, 33, 34]), and C++ standard SIMD features that have been added to C++26 [35] to help ensure full code portability of our application. Here, we used an early version of the SIMD features (std::experimental::simd) to allow us to stay compatible with older C++ versions (Octo-Tiger can still be compiled using C++20 compilers). The use of Kokkos, HPX, and SIMD Features guarantees its seamless portability and consistent scaling characteristics across a wide variety of heterogeneous (GPU and non-GPU based) hardware platforms, in particular on Intel, AMD, and NVIDIA products, but also on upcoming ARM64 and RISC-V HPC architectures.
In the problem domain investigated, the amount of computational work per compute kernel is low, requiring adaptive work aggregation on GPU devices to reduce scheduling overheads and starvation effects for the related asynchronous tasks [2]. The nature of highly complex 3D AMR applications like Octo-Tiger (see Figure 2) require exascale computing systems to solve underlying physics problems that require very large amounts of data and pose computational challenges. To achieve high parallel efficiency, these challenges include providing intra- and inter-node load-balancing, pervasive overlapping of computation and communication, tight integration and high utilization of accelerator devices, and runtime-adaptive task placement and scheduling across nodes and diverse execution environments.
@@ 4.1 HPX
HPX is an asynchronous many-task runtime system (AMT). It allows users to express the parallelism within their application by dynamically building a task-graph on-the-fly using HPX futures and continuations. For example, a function can be launched asynchronously using hpx::async, which returns an HPX future. This future can be used to suspend the current HPX task (get) or define another task that will be automatically triggered once the original function is done (then). Each task can subsequently spawn an arbitrary number of other tasks asynchronously and with little overhead, making parallel work available quickly, which is especially useful for parallelized tree traversals. Of course, there is a lot more functionality available. Particularly, HPX implements all C++23 APIs regarding parallelism and concurrency. For a full list of the functionality available, we refer to HPX’s online documentation [^1] https://hpx-docs.stellar-group.org/latest/html/index.html. The tasks within HPX’s task graph are then processed by a set of HPX worker threads. Thus, HPX can easily handle billions of tasks, which are processed by the worker threads as they become available within the graph.
@@
HPX also contains various key features to enable the development of distributed codes [36]. For example, functions can also be launched asynchronously on other compute-nodes, still getting an HPX future back. There are also communication channels and other distributed features available that tie into the HPX task graph similarly. This allows users to build a distributed task graph. This is made possible by HPX’s Active Global Address Space (AGAS) and various networking backends (parcelports).
@@
Overall, HPX is well suited to address the computational challenges of developing scalable, distributed AMR applications, directly addressing some of the challenges mentioned in Section The computational challenges. With HPX, users can exploit the parallelism within tree-structures more easily by expressing it in a similar graph structure, the HPX task-graph. Through this task graph, which may contain both computation and communication tasks, users can automatically achieve an overlapping of computation and communication. This helps one to scale their codes to numerous compute nodes with a high parallel efficiency. Furthermore, thanks to some of our integrations, we can also incorporate accelerator devices into this task-graph. This helps to keep a high utilization of the accelerator devices by better overlapping their computations with the CPU tasks and communication.
@@
Use of HPX can also help with portability. HPX’s portability is achieved through an interface (API) fully aligned with the latest C++ standard specification. The API decouples the application from the upper layers of the runtime system and hides the low-level details of the underlying architecture. HPX provides general-purpose building blocks (such as C++ standards conforming parallel algorithms), as well as high-level utilities coordinating asynchronous execution, and offers a programming abstraction that facilitates the programming of both shared-memory and distributed-memory systems through a uniform programming interface, scaling up from hand-held devices to HPC clusters.
@@
In terms of networking portability, HPX’s modular architecture abstracts the concrete networking hardware allowing for Octo-Tiger to be completely independent of it. HPX supports TCP/IP, MPI, LCI [6, 37], OpenSHMEM [38], and GASNet [39] as possible networking conduits. In this work, we focus on two backends (parcelports): MPI and LCI.